# TC-KANRecon: High-Quality and Accelerated MRI Reconstruction via Adaptive KAN Mechanisms and Intelligent Feature Scaling

### v1

@@ Abstract
Magnetic Resonance Imaging (MRI) has become essential in clinical diagnosis due to its high resolution and multiple contrast mechanisms. However, the relatively long acquisition time limits its broader application. To address this issue, this study presents an innovative conditional guided diffusion model, named as TC-KANRecon, which incorporates the Multi-Free U-KAN (MF-UKAN) module and a dynamic clipping strategy. TC-KANRecon model aims to accelerate the MRI reconstruction process through deep learning methods while maintaining the quality of the reconstructed images. The MF-UKAN module can effectively balance the tradeoff between image denoising and structure preservation. Specifically, it presents the multi-head attention mechanisms and scalar modulation factors, which significantly enhances the model’s robustness and structure preservation capabilities in complex noise environments. Moreover, the dynamic clipping strategy in TC-KANRecon adjusts the cropping interval according to the sampling steps, thereby mitigating image detail loss typically caused by traditional cropping methods and enriching the visual features of the images. Furthermore, the MC-Model module incorporates full-sampling k-space information, realizing efficient fusion of conditional information, enhancing the model’s ability to process complex data, and improving the realism and detail richness of reconstructed images. Experimental results demonstrate that the proposed method outperforms other MRI reconstruction methods in both qualitative and quantitative evaluations. Notably, TC-KANRecon method exhibits excellent reconstruction results when processing high-noise, low-sampling-rate MRI data. Our source code is available at https://github.com/lcbkmm/TC-KANRecon.
@@ 1 Introduction
where and denote the fully-sampled and undersampled k-space signal, respectively, represents the undersampling mask, and is the Fourier operator. The goal is to find a MR image such that its k-space content aligns with , often referred to as the data consistency term. Furthermore, should adhere to certain prior knowledge about MR images, as expressed by the regularization term R(), which is subject to numerous innovations.
@@
Although undersampling reduces the amount of acquired data and shortens scan times, it violates the Nyquist-Shannon sampling theorem, potentially introducing aliasing artifacts. To address this, researchers have proposed Compressed Sensing (CS) technology [4] to mitigate aliasing artifacts caused by undersampling. CS formulates image reconstruction as an optimization problem, incorporating assumptions of sparsity and incoherence. However, the complexity and limitations of CS have hindered its widespread application due to the need for multiple iterative computations and substantial computational resources. Additionally, the effectiveness of CS methods heavily relies on empirical hyperparameter selection, necessitating individualized adjustments under different scanning conditions, a time-consuming and laborintensive process [5]. To further improve MRI acceleration capabilities, researchers have explored various methods such as low-rank constraints [6], adaptive sparse modeling [7], and parallel imaging techniques [8]. These methods enhance image reconstruction quality and speed through diverse technical approaches but still face challenges with aliasing artifacts at high acceleration factors [9]. Parallel imaging techniques, which use multiple receiver coils to simultaneously acquire signals, can enhance acquisition speed and image quality. However, their effectiveness depends on coil configuration and imaging parameter selection, requiring experienced operators for tuning.
@@
In recent years, the application of artificial intelligence technology in the field of MRI has infused new vitality into MRI acceleration and reconstruction techniques. Through training, deep learning models are capable of reconstructing high-quality images from undersampled k-space data, drastically simplifying the complex parameter tuning processes of traditional methods. Their efficient computational capabilities further drastically reduce image reconstruction time, enhancing the efficiency of medical services. However, despite breakthroughs in existing deep learning technologies such as diffusion models, challenges persist. Noise sensitivity is particularly problematic, as device noise, patient movement, and other factors in MRI data can easily interfere with image quality, leading models to mistakenly learn noise, resulting in artifacts or blurred images that compromise diagnostic accuracy. Additionally, inadequate preservation of structural information poses a significant challenge. The fine anatomical structures in MR images are crucial for disease diagnosis, yet existing methods often struggle to retain these details while achieving rapid reconstruction. Furthermore, the lack of robustness in deep learning models limits their clinical application. The significant variations in MRI data among different patients, scanning protocols, and devices require models to possess strong generalization capabilities. However, current models are often trained on specific datasets, and their adaptability to new environments needs improvement. To address these challenges, we innovatively propose the MF-UKAN (Multi-Feature U-KAN) module and its complementary dynamic cropping strategy, significantly enhancing the precision, efficiency, and flexibility of diffusion models in MR image reconstruction tasks. Our method offers several key contributions and advantages:
1.
@@
To balance denoising and structural information preservation, we designed the MF-UKAN module. This design incorporates multi-head attention mechanisms and scalar modulation factors for fine-grained control over backbone and skip features, enhancing the robustness of model in complex noise environments and improving the preservation of structural information in reconstructed images.
2.
@@
We introduced an innovative dynamic clipping strategy to address the limitations of traditional cropping methods, which restrict image diversity. Our strategy dynamically adjusts cropping interval boundaries based on sampling steps, effectively mitigating image detail loss and resulting in images with richer visual features and reasonable brightness distribution.
3.
@@
We combined the fully sampled k-space information with MRI data in the MC-Model module and processed it using the encoder stage of the MF-UKAN module. This approach enhances the model’s ability to handle complex data and facilitates targeted MR images generation, improving the realism and detail richness of reconstructed images.
4.
@@
We conducted extensive comparative experiments on two public datasets, and the results indicated that our method surpassed other state-of-the-art MRI reconstruction methods. Specifically, it demonstrates superior performance and faster inference speed when processing high-noise, low-sampling-rate MRI data. Ablation experiments further validated the critical role of each module in the TC-KANRecon model, proving the effectiveness and necessity of its design.
@@ 2 Related work
With the rapid advancement of deep learning, models leveraging this technology have become increasingly prominent in MRI reconstruction, demonstrating exceptional performance. Early approaches [10] utilized single feedforward Convolutional Neural Networks (CNNs), such as SRCNN [11] and U-net [12], to map undersampled k-space data to fully sampled images in an end-to-end manner. More sophisticated models have since emerged, adopting iterative architectures that break down the reconstruction process into several learnable optimization stages. For instance, Sun et al. [13] introduced ADMM-Net for MRI reconstruction, which represents each stage as an iteration of the Alternating Direction Method of Multipliers (ADMM) algorithm [14]. Similarly, rsGAN [15] employs conditional generative adversarial networks for the joint recovery of undersampled multi-contrast acquisitions. To emulate iterative dictionary learning methods, Schlemper et al. [16] and Dar et al. [17] proposed deep cascaded CNN architectures, which perform convolution operations in image space through multiple residual learning blocks, Zeng et al. [18] applied CNN in MRI reconstruction. Qin et al. [19] further advanced this concept by introducing a Convolutional Recurrent Neural Network (CRNN) architecture that leverages time-series dependencies and iterative optimization benefits. Subsequent improvements to CRNN designs by Wang et al. [20], Chen et al. [21], and Guo et al. [22] incorporated recursive pyramid layers, neural ODEs, and overcomplete network architectures in the hope of further enhancing reconstruction quality through the above strategies.
@@
Despite the ability of GAN-based methods to generate realistic image samples, their limited representational diversity can hinder reconstruction performance. Recent developments in diffusion models have shown potential to address this issue. These models transform Gaussian noise into image samples through a multi-step process, directly characterizing data distribution correlations. By combining learned priors with imaging operators during inference, diffusion models enable repeated projection reconstruction. For instance, Jalal et al. [28] and Luo et al. [29] proposed using score functions and Langevin dynamics for sampling. Chung et al. [30] and Song et al. [31] employed predictors to solve stochastic differential equations. Peng et al. [32] introduced DiffuseRecon, a diffusion model-based MRI reconstruction method that does not require additional training for specific acceleration factors. Cao et al. [33] proposed a complex diffusion probabilistic model that better preserves the complex-valued information of MRI data. Gungor et al. [34] developed Adaptive Diffusion Priors, such as AdaDiff, to enhance reconstruction performance during inference. Cao et al. [35] also designed a high-frequency DDPM model to retain high-frequency information in MRI data. However, these diffusion-based methods have not fully addressed the inherent limitations of the U-net architecture used in diffusion models, resulting in overly smooth images and suboptimal imaging quality.
@@ 3.1 MF-UKAN Module
Inspired by Li et al. [38], we substitute the U-net architecture in stable diffusion [39] with a U-KAN network. The Tok-KAN module consists of a tokenization layer, a KAN layer, and a Batch Normalization (BN) layer. The structure of the Tok-KAN module is depicted in Fig. 1.
@@
where is the output feature map of the -th layer, is a linear projection, and denotes the temporal embedding for a given time step.
The U-KAN, characterized by its U-net-like structure, facilitates the denoising process through the backbone path and transmits high-frequency detailed features to the decoder part with the help of skip connections. However, this design can overemphasize high-frequency information, potentially weakening the backbone network’s ability to capture essential semantic features for denoising. To address this, we propose an improved U-KAN module, MF-UKAN, which introduces a multi-attention mechanism and two critical scalar modulation factors. The backbone feature scaling factor () aims to enhance the expressive power of the backbone feature map (), while the skip feature scaling factor () moderately attenuates the influence of the skip feature map () to avoid excessive high-frequency interference.
@@ 3.2 MC-Model
To enhance the directional generation of MR images, we use fully sampled k-space information as a conditional guide. Initially, we apply an inverse Fourier transform to the fully sampled k-space data, denoted as , and subsequently fuse it with the MR images through concatenation. This process is represented by Eq.(10):
@@
To maintain consistency with the diffusion backbone’s input, which is potential space, we compress using an additional convolution module. This module employs a kernel size of 3, a step size of 1, and a padding of 1. The compression process involves an initial convolutional layer, followed by a ReLU activation function and a group normalization layer. This sequence is executed three times to produce a feature map with the same shape as the diffusion backbone, which is then passed to subsequent modules. The final feature vector is described by Eq.(11):
@@ 3.3 Dynamic Clipping Strategy
Recently, Chen et al. [40] proposed an innovative dynamic clipping strategy to replace the conventional fixed clipping approach, which strictly confines the predicted variable within the interval . This new strategy involves dynamically adjusting the clipping interval boundaries based on the current sampling step , defining a varying interval , where evolves with the sampling process. They formulated a linear model as shown in Eq.(12):
This strategy allows for flexible adjustment of the clipping threshold according to the model’s training progress and the stage of image generation. It provides greater freedom during the initial stages of image structure formation, resulting in richer visual features. However, during the early stages of diffusion model sampling, when the model generates blurry, low-resolution initial image structures from pure noise, a larger clipping threshold is needed to increase freedom and reduce information loss, aiding in capturing overall shapes. As sampling progresses and image resolution and details increase, the clipping threshold should gradually decrease to limit noise and enhance image finesse and quality. Consequently, we have refined the linear model to Eq.(13):
This revised model allows the clipping threshold to linearly decrease as the sampling step increases. This dynamic clipping strategy preserves image diversity during the initial stages of diffusion model sampling and retains image details and precision as sampling deepens. Consequently, it produces images that better align with real-world counterparts, as illustrated in Fig. 2.
@@ 4.3 Evaluation Metrics
To thoroughly assess the performance of our model, we employed three evaluation metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Normalized Mean Square Error (NMSE). PSNR quantifies the ratio of peak signal energy to average noise energy, offering a reliable measure of the differences between images before and after reconstruction. SSIM assesses the similarity between images by evaluating brightness, contrast, and structure, thereby indicating how visually similar the reconstructed image is to the original. A higher SSIM value signifies a closer resemblance to the original image NMSE normalizes the mean square error by dividing it by the mean square error of the original image. A lower NMSE value indicates a closer resemblance to the original image and therefore higher image quality. Utilizing these three metrics allows for a comprehensive and accurate evaluation of our MRI reconstruction method. This approach fully captures the differences and similarities between the reconstructed and original images, thereby ensuring the reliability of the model’s performance.
@@ 4.4 Comparison Experiments
As shown in Tables 1 and 2, we conducted comparative experiments with other state-of-the-art models for MRI reconstruction on the fastMRI and SKM-TEA datasets to validate the efficacy of the TC-KANRecon model. The experimental results indicate that the TC-KANRecon model significantly outperforms its counterparts in terms of PSNR, SSIM, and NMSE metrics.
@@
The innovative network structure and diversified feature processing strategies of the TC-KANRecon model enhance its adaptability and robustness in MRI reconstruction tasks. Consequently, it surpasses existing models in all evaluated metrics, demonstrating its superior performance. The comparative results with other state-of-the-art models, specifically for an Acceleration Factor (AF) of 4, are illustrated in Fig. 3.
Given the consistent results across different single-coil datasets and various acceleration factors, we focus our detailed analysis on the reconstruction performance of the model using the fastMRI dataset with AF=4, both in comparison and ablation experiments.
Compared to the U-net model, although U-net performs well in many image processing tasks, it struggles with capturing high-frequency details and dealing with noise in MR images due to its encoder-decoder structure, which can result in the loss of detail during feature extraction. For instance, the PSNR of the U-net model is 27.68, while the TC-KANRecon model achieves 30.50. Similarly, the SSIM for U-net is 0.631, significantly lower than TC-KANRecon’s 0.817. In terms of NMSE, U-net scores 0.0295, whereas TC-KANRecon registers a lower 0.0211. The TC-KANRecon model, incorporating the MF Model, optimizes the balance of skip connections and main features, substantially enhancing reconstruction accuracy.
@@
Compared to the KIKI-net model, which struggles with high-frequency detail processing and denoising due to its lack of effective strategies for capturing fine details and suppressing noise, the TC-KANRecon model demonstrates superior performance. The innovative texture coordination strategy and multi-attention mechanism of TC-KANRecon enhance detail restoration helping the model excel in restoring high-frequency details and effectively managing noise, significantly improving overall image quality. Specifically, when compared with KIKI-net, TC-KANRecon increases the PSNR from 27.01 to 30.50 and the SSIM from 0.613 to 0.817. In terms of NMSE metrics, KIKI-net scores 0.0354, while TC-DiffRecon achieves a lower value of 0.0211.
@@
Compared to the D5C5 model, which shows some advantages in MRI reconstruction, D5C5’s static feature extraction methods result in partial detail loss, affecting detail restoration and overall image quality. The PSNR for D5C5 is 27.74, while TC-KANRecon achieves 30.50. The SSIM for D5C5 is 0.632, lower than TC-KANRecon’s 0.817. NMSE metrics for U-net are 0.0289, while TC-KANRecon scores 0.0211. TC-KANRecon, with its multi-attention mechanism and dynamic clipping strategy, significantly outperforms D5C5 in reducing details and overall image quality, demonstrating superior image detail restoration capability.
@@
Compared to the ReconFormer model, which excels in capturing long-range dependencies using a Transformer structure, TC-KANRecon demonstrates improved performance in handling complex details of MR images. While ReconFormer effectively captures global information, it often overlooks local detail features, leading to suboptimal performance. In contrast, TC-KANRecon combines the MF Model and the Tok-KAN module to enhance feature extraction and model interpretability. This results in superior performance in metrics such as PSNR and SSIM. Specifically, TC-KANRecon achieves a PSNR of 30.50, compared to ReconFormer’s 28.62. Additionally, the SSIM for TC-KANRecon reaches 0.817, while ReconFormer attains 0.643. Regarding NMSE metrics, ReconFormer scores 0.0263, whereas TC-KANRecon achieves a lower value of 0.0211.

### v2

@@ Abstract
Magnetic Resonance Imaging (MRI) has become essential in clinical diagnosis due to its high resolution and multiple contrast mechanisms. However, the relatively long acquisition time limits its broader application. To address this issue, this study presents an innovative conditional guided diffusion model, named as TC-KANRecon, which incorporates the Multi-Free U-KAN (MF-UKAN) module and a dynamic clipping strategy. TC-KANRecon model aims to accelerate the MRI reconstruction process through deep learning methods while maintaining the quality of the reconstructed images. The MF-UKAN module can effectively balance the tradeoff between image denoising and structure preservation. Specifically, it presents the multi-head attention mechanisms and scalar modulation factors, which significantly enhances the model’s robustness and structure preservation capabilities in complex noise environments. Moreover, the dynamic clipping strategy in TC-KANRecon adjusts the cropping interval according to the sampling steps, thereby mitigating image detail loss typicalching the visual features of the images. Furthermore, the MC-Model incorporates full-sampling k-space information, realizing efficient fusion of conditional information, enhancing the model’s ability to process complex data, and improving the realism and detail richness of reconstructed images. Experimental results demonstrate that the proposed method outperforms other MRI reconstruction methods in both qualitative and quantitative evaluations. Notably, TC-KANRecon method exhibits excellent reconstruction results when processing high-noise, low-sampling-rate MRI data. Our source code is available at https://github.com/lcbkmm/TC-KANRecon.
@@ 1 Introduction
where and denote the fully-sampled and undersampled k-space signal, respectively, represents the undersampling mask, and is the Fourier operator. The goal is to find a MR image such that its k-space content aligns with , often referred to as the data consistency term. Furthermore, should adhere to certain prior knowledge about MR images, as expressed by the regularization term , which is subject to numerous innovations.
@@
Although undersampling reduces the amount of acquired data and shortens scan times, it violates the Nyquist-Shannon sampling theorem, potentially introducing aliasing artifacts. To address this, researchers have proposed Compressed Sensing (CS) technology [6] to mitigate aliasing artifacts caused by undersampling. CS formulates image reconstruction as an optimization problem, incorporating assumptions of sparsity and incoherence. However, the complexity and limitations of CS have hindered its widespread application due to the need for multiple iterative computations and substantial computational resources. Additionally, the effectiveness of CS methods heavily relies on empirical hyperparameter selection, necessitating individualized adjustments under different scanning conditions, a time-consuming and laborintensive process [7]. To further improve MRI acceleration capabilities, researchers have explored various methods such as low-rank constraints [8], adaptive sparse modeling [9], applying the TVJ1-ESPIRiT method in combination with PI and CS [10] and parallel imaging techniques [11]. These methods enhance image reconstruction quality and speed through diverse technical approaches but still face challenges with aliasing artifacts at high acceleration factors [12]. Parallel imaging techniques, which use multiple receiver coils to simultaneously acquire signals, can enhance acquisition speed and image quality. However, their effectiveness depends on coil configuration and imaging parameter selection, requiring experienced operators.
@@
In recent years, the application of artificial intelligence technology in the field of MRI has infused new vitality into MRI acceleration and reconstruction techniques. Through training, deep learning models are capable of reconstructing high-quality images from undersampled k-space data, drastically simplifying the complex parameter tuning processes of traditional methods. Their efficient computational capabilities further drastically reduce image reconstruction time, enhancing the efficiency of medical services. However, despite breakthroughs in existing deep learning technologies such as diffusion models, challenges persist. Noise sensitivity is particularly problematic, as device noise, patient movement, and other factors in MRI data can easily interfere with image quality, leading models to mistakenly learn noise, resulting in artifacts or blurred images that compromise diagnostic accuracy. Additionally, inadequate preservation of structural information poses a significant challenge. The fine anatomical structures in MR images are crucial for disease diagnosis, yet existing methods often struggle to retain these details while achieving rapid reconstruction. Furthermore, the lack of robustness in deep learning models limits their clinical application. The significant variations in MRI data among different patients, scanning protocols, and devices require models to possess strong generalization capabilities. However, current models are often trained on specific datasets, and their adaptability to new environments needs improvement. To address these challenges, we innovatively propose the Multi-Free U-KAN (MF-UKAN) module and its complementary dynamic cropping strategy, significantly enhancing the precision, efficiency, and flexibility of diffusion models in MR image reconstruction tasks. Our method offers several key contributions and advantages:
@@
To balance denoising with the preservation of structural information, we introduced the MF-UKAN module. This module employs multi-head attention mechanisms and scalar modulation factors, providing fine-grained control over backbone and skip features. This design greatly enhances the model’s robustness in noisy environments and also improves the retention of structural details in reconstructed images.
@@
We introduced an innovative dynamic clipping strategy to overcome the limitations of traditional cropping methods, which often restrict image diversity. Our strategy dynamically adjusts cropping interval boundaries based on sampling steps, effectively reducing image detail loss. This results in images with richer visual features and better brightness distribution.
@@
We integrated fully sampled k-space information with MRI data in the MC-Model module, processing it through the encoder stage of the MF-UKAN module. This combination enhances the model’s capability to handle complex data and aids in generating targeted MR images, thus improving the realism and detail in reconstructed images.
@@
We conducted extensive comparative experiments on two public datasets. Our results indicate that our method outperforms other state-of-the-art MRI reconstruction methods, particularly in terms of performance when processing high-noise, low-sampling-rate MRI data. Ablation experiments further validated the critical role of each module in the TC-KANRecon model, confirming the effectiveness and necessity of our design.
@@ 2 Related work
With the rapid advancement of deep learning, models leveraging this technology have become increasingly prominent in MRI reconstruction, demonstrating exceptional performance. Early approaches [13] utilized single feedforward Convolutional Neural Networks (CNNs), such as SRCNN [14] and U-Net [15], to map undersampled k-space data to fully sampled images in an end-to-end manner. More sophisticated models have since emerged, adopting iterative architectures that break down the reconstruction process into several learnable optimization stages. For instance, Sun et al. [16] introduced ADMM-Net for MRI reconstruction, which represents each stage as an iteration of the Alternating Direction Method of Multipliers (ADMM) algorithm [17]. Similarly, Dar et al. [18] employs conditional generative adversarial networks for the joint recovery of undersampled multi-contrast acquisitions. To emulate iterative dictionary learning methods, Schlemper et al. [19] and Dar et al. [20] proposed deep cascaded CNN architectures, which perform convolution operations in image space through multiple residual learning blocks, Zeng et al. [21] applied CNN in MRI reconstruction. Qin et al. [22] further advanced this concept by introducing a Convolutional Recurrent Neural Network (CRNN) architecture that leverages time-series dependencies and iterative optimization benefits. Subsequent improvements to CRNN designs by Wang et al. [23], Chen et al. [24], and Guo et al. [25] incorporated recursive pyramid layers, neural ODEs, and overcomplete network architectures in the hope of further enhancing reconstruction quality through the above strategies. In addition, Machado et al. [26] proposes method adds the undersampling factor in the reconstruction process and combining it with other tasks, achieving faster and better reconstruction results.
@@
While GAN-based methods are renowned for their ability to generate realistic images, they often suffer from a lack of diverse representation, which can impede their performance in image reconstruction. Recent advancements in diffusion models offer promising solutions to this limitation. These models convert Gaussian noise into image samples through a multi-step procedure that directly maps the correlations within the data distribution. By integrating learned priors with imaging operators during the inference stage, diffusion models enable repeated projections for improved reconstruction. For example, Jalal et al. [32] and Luo et al. [33] introduced the use of score functions and Langevin dynamics for sampling. Similarly, Song et al. [34] utilized predictors to address stochastic differential equations in the methodology. Peng et al. [35] developed DiffuseRecon, a diffusion model-based MRI reconstruction approach that eliminates the need for additional training for different acceleration factors. Meanwhile, the CDPM model proposed by Cao et al. [36] effectively preserves the complex-valued information critical in MRI data. Gungor et al. [37] created Adaptive Diffusion Priors, such as AdaDiff, to further enhance reconstruction performance during inference. Cao et al. [38] also designed a high-frequency DDPM model to retain high-frequency information in MRI data. Despite these advancements, diffusion-based methods have not completely overcome the inherent limitations of the U-Net architecture employed within these models. This shortcoming often results in overly smooth images and suboptimal imaging quality, indicating areas for future enhancement and research.
@@ 3.1 Multi-Free U-KAN
Inspired by U-KAN [41], we substitute the U-Net architecture in stable diffusion [42] with a U-KAN network. The Tok-KAN module consists of a tokenization layer, a KAN layer, and a Batch Normalization (BN) layer. The structure of the Tok-KAN module is depicted in Fig. 1.
@@
where is the output feature map of the -th layer, F is a linear projection, and denotes the temporal embedding for a given time step.
The U-KAN, characterized by its U-Net-like structure, facilitates the denoising process through the backbone path and transmits high-frequency detailed features to the decoder part with the help of skip connections. However, this design can overemphasize high-frequency information, potentially weakening the backbone network’s ability to capture essential semantic features for denoising. To address this issue, we propose an improved U-KAN module, MF-UKAN, inspired by the MF Model proposed by Zhang et al. [43], which introduces a multi-attention mechanism and two critical scalar modulation factors. The backbone feature scaling factor aims to enhance the expressive power of the backbone feature map , while the skip feature scaling factor moderately attenuates the influence of the skip feature map to avoid excessive high-frequency interference.
@@ 3.2 MC-Model
To enhance the directional generation of MR images, we use fully sampled k-space information as a conditional guide. Initially, we apply an inverse Fourier transform to the fully sampled k-space data, denoted as , and subsequently fuse it with the MR images through concatenation. This process is represented by Eq. 10:
@@
To maintain consistency with the diffusion backbone’s input, which is potential space, we compress using an additional convolution module. This module employs a kernel size of 3, a step size of 1, and a padding of 1. The compression process involves an initial convolutional layer, followed by a ReLU activation function and a group normalization layer. This sequence is executed three times to produce a feature map with the same shape as the diffusion backbone, which is then passed to subsequent modules. The final feature vector is described by Eq. 11:
@@ 3.3 Dynamic Clipping Strategy
Recently, Sun et al. [44] proposed an innovative temporal adaptive thresholding strategy to replace the conventional fixed clipping approach, which strictly confines the predicted variable within the interval . This new strategy involves dynamically adjusting the clipping interval boundaries based on the current sampling step , defining a varying interval , where evolves with the sampling process. They formulated a linear model as shown in Eq. 12:
This strategy allows for flexible adjustment of the clipping threshold according to the model’s training progress and the stage of image generation. It provides greater freedom during the initial stages of image structure formation, resulting in richer visual features. However, during the early stages of diffusion model sampling, when the model generates blurry, low-resolution initial image structures from pure noise, a larger clipping threshold is needed to increase freedom and reduce information loss, aiding in capturing overall shapes. As sampling progresses and image resolution and details increase, the clipping threshold should gradually decrease to limit noise and enhance image finesse and quality. Consequently, we have refined the linear model to Eq. 13:
This revised model allows the clipping threshold to linearly decrease as the sampling step increases. This dynamic clipping strategy preserves image diversity during the initial stages of diffusion model sampling and retains image details and precision as sampling deepens. Consequently, it produces images that better align with real-world counterparts, as illustrated in Fig. 2
@@ 4.3 Evaluation Metrics
To thoroughly assess the performance of our model, we employed three evaluation metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Normalized Mean Square Error (NMSE). PSNR measures the peak signal to noise ratio, indicating image quality by comparing signal energy to noise. SSIM compares image similarity in terms of brightness, contrast, and structure, with higher values meaning the reconstructed image looks more like the original. NMSE normalizes the mean square error against the original image’s error, with lower values indicating higher quality. These metrics provide a comprehensive assessment of our model’s performance in reconstructing MRI images accurately.
@@ 4.4 Comparison Experiments
As shown in Tables 1 and 2, we conducted comparative experiments with Zero Filled (ZF) and other state-of-the-art models for MRI reconstruction on the fastMRI and SKM-TEA datasets to validate the efficacy of the TC-KANRecon model. The experimental results indicate that the TC-KANRecon model significantly outperforms its counterparts in terms of PSNR, SSIM, and NMSE metrics. The innovative network structure and diversified feature processing strategies of the TC-KANRecon model enhance its adaptability and robustness in MRI reconstruction tasks. Consequently, it surpasses existing models in all evaluated metrics, demonstrating its superior performance. The comparative results with other state-of-the-art models, specifically for an Acceleration Factor (AF) of 4, are illustrated in Fig. 3. Given the consistent results across different single-coil datasets and various acceleration factors, we focus our detailed analysis on the reconstruction performance of the model using the fastMRI dataset with AF = 4, both in comparison and ablation experiments.
@@
Compared to the KIKI-net model [47], which struggles with high-frequency detail processing and denoising due to its lack of effective strategies for capturing fine details and suppressing noise, the TC-KANRecon model demonstrates superior performance. The innovative texture coordination strategy and multi-attention mechanism of TC-KANRecon enhance detail restoration helping the model excel in restoring high-frequency details and effectively managing noise, significantly improving overall image quality. Specifically, when compared with KIKI-net, TC-KANRecon increases the PSNR from 27.01 to 30.50 and the SSIM from 0.613 to 0.817. In terms of NMSE metrics, KIKI-net scores 0.0354, while TC-DiffRecon achieves a lower value of 0.0211.
@@
Compared to the D5C5 model [19], which shows some advantages in MRI reconstruction, D5C5’s static feature extraction methods result in partial detail loss, affecting detail restoration and overall image quality. The PSNR for D5C5 is 27.74, while TC-KANRecon achieves 30.50. The SSIM for D5C5 is 0.632, lower than TC-KANRecon’s 0.817. NMSE metrics for U-Net are 0.0289, while TC-KANRecon scores 0.0211. TC-KANRecon, with its multi-attention mechanism and dynamic clipping strategy, significantly outperforms D5C5 in reducing details and overall image quality, demonstrating superior image detail restoration capability.
Compared to the U-Net model [15], although U-Net performs well in many image processing tasks, it struggles with capturing high-frequency details and dealing with noise in MR images due to its encoder-decoder structure, which can result in the loss of detail during feature extraction. For instance, the PSNR of the U-Net model is 27.68, while the TC-KANRecon model achieves 30.50. Similarly, the SSIM for U-Net is 0.631, significantly lower than TC-KANRecon’s 0.817. In terms of NMSE, U-Net scores 0.0295, whereas TC-KANRecon registers a lower 0.0211. The TC-KANRecon model, incorporating the MF Model, optimizes the balance of skip connections and main features, substantially enhancing reconstruction accuracy.
Compared to the OUCR model [25], TC-KANRecon demonstrates significant advantages in MRI reconstruction. While the OUCR model captures both local and global features by combining overcomplete and undercomplete branches, it falls short in detail restoration and overall image quality. Specifically, the OUCR model achieves a PSNR of 28.08, whereas TC-KANRecon attains 30.05. The SSIM for the OUCR model is 0.633, markedly lower than TC-KANRecon’s 0.817. Additionally, the NMSE metric for the OUCR model is 0.0277, compared to TC-KANRecon’s improved performance of 0.0211. TC-KANRecon leverages an innovative conditional guided diffusion model, which integrates the MF-UKAN module and a dynamic clipping strategy, effectively balancing image denoising and structure preservation.
fastMRI [45] | ||||||||||||
Method | PSNR() | SSIM() | NMSE() | |||||||||
AF Factor | 4 | 6 | 8 | 10 | 4 | 6 | 8 | 10 | 4 | 6 | 8 | 10 |
ZF | 22.91 | 20.79 | 18.37 | 15.07 | 0.549 | 0.482 | 0.414 | 0.349 | 0.0436 | 0.0488 | 0.0567 | 0.0719 |
KIKI-net [47] | 27.01 | 26.55 | 25.08 | 24.06 | 0.613 | 0.503 | 0.492 | 0.412 | 0.0354 | 0.0397 | 0.0462 | 0.0542 |
D5C5 [19] | 27.74 | 26.56 | 26.12 | 24.60 | 0.632 | 0.529 | 0.512 | 0.445 | 0.0289 | 0.0374 | 0.0440 | 0.0532 |
U-Net [15] | 27.68 | 26.23 | 25.83 | 24.49 | 0.631 | 0.520 | 0.511 | 0.436 | 0.0295 | 0.0382 | 0.0430 | 0.0520 |
OUCR [25] | 28.08 | 27.66 | 26.23 | 25.31 | 0.633 | 0.534 | 0.516 | 0.482 | 0.0277 | 0.0368 | 0.0389 | 0.0486 |
DiffuseRecon [35] | 28.27 | 27.83 | 26.60 | 25.90 | 0.635 | 0.586 | 0.549 | 0.528 | 0.0272 | 0.0360 | 0.0381 | 0.0466 |
ReconFormer [48] | 28.62 | 28.01 | 26.64 | 26.08 | 0.643 | 0.590 | 0.529 | 0.519 | 0.0263 | 0.0348 | 0.0374 | 0.0480 |
TC-DiffRecon [43] | 29.78 | 28.77 | 27.35 | 26.49 | 0.742 | 0.637 | 0.591 | 0.562 | 0.0247 | 0.0313 | 0.0338 | 0.0415 |
TC-KANRecon (Ours) | 30.50 | 29.16 | 27.93 | 26.83 | 0.817 | 0.691 | 0.624 | 0.593 | 0.0211 | 0.0273 | 0.0309 | 0.0371 |
SKM-TEA [46] | ||||||||||||
Method | PSNR() | SSIM() | NMSE() | |||||||||
AF Factor | 4 | 6 | 8 | 10 | 4 | 6 | 8 | 10 | 4 | 6 | 8 | 10 |
ZF | 23.63 | 22.09 | 20.06 | 16.72 | 0.653 | 0.587 | 0.529 | 0.455 | 0.0361 | 0.0422 | 0.0511 | 0.0641 |
KIKI-net [47] | 28.16 | 26.96 | 25.80 | 25.11 | 0.716 | 0.620 | 0.601 | 0.543 | 0.0286 | 0.0350 | 0.0434 | 0.0510 |
D5C5 [19] | 28.87 | 28.24 | 27.35 | 27.01 | 0.731 | 0.654 | 0.637 | 0.586 | 0.0221 | 0.0260 | 0.0384 | 0.0460 |
U-Net [15] | 28.52 | 27.79 | 26.73 | 26.34 | 0.723 | 0.645 | 0.631 | 0.571 | 0.0254 | 0.0336 | 0.0410 | 0.0493 |
OUCR [25] | 29.26 | 28.90 | 28.06 | 27.24 | 0.745 | 0.656 | 0.642 | 0.603 | 0.0165 | 0.0247 | 0.0320 | 0.0387 |
DiffuseRecon [35] | 29.51 | 29.13 | 28.20 | 27.49 | 0.749 | 0.712 | 0.682 | 0.668 | 0.0152 | 0.0232 | 0.0303 | 0.0352 |
ReconFormer [48] | 29.74 | 29.43 | 28.77 | 28.37 | 0.765 | 0.724 | 0.693 | 0.654 | 0.0149 | 0.0216 | 0.0278 | 0.0326 |
TC-DiffRecon [43] | 30.59 | 30.26 | 29.50 | 29.22 | 0.852 | 0.771 | 0.732 | 0.708 | 0.0136 | 0.0198 | 0.0254 | 0.0311 |
TC-KANRecon (Ours) | 31.48 | 30.93 | 30.12 | 29.74 | 0.893 | 0.822 | 0.785 | 0.741 | 0.0118 | 0.0189 | 0.0242 | 0.0274 |
@@
Compared to the ReconFormer model [48], which excels in capturing long-range dependencies using a Transformer structure, TC-KANRecon demonstrates improved performance in handling complex details of MR images. While ReconFormer effectively captures global information, it often overlooks local detail features, leading to suboptimal performance. In contrast, TC-KANRecon combines the MF Model and the Tok-KAN module to enhance feature extraction and model interpretability. This results in superior performance in metrics such as PSNR and SSIM. Specifically, TC-KANRecon achieves a PSNR of 30.50, compared to ReconFormer’s 28.62. Additionally, the SSIM for TC-KANRecon reaches 0.817, while ReconFormer attains 0.643. Regarding NMSE metrics, ReconFormer scores 0.0263, whereas TC-KANRecon achieves a lower value of 0.0211.